# Seungone Kim1,2,3  Juyoung $\textbf{}^{*}$  Shayne Longpre4Bill Yuchen Lin5Jamin Shin1 Sean Welleck3 Graham Neubig3Moontae Lee2,6Kyungjae Lee2Minjoon Seo1

### v1

@@ 6.1 Weight Merging vs Joint Training
Surprisingly, we observe that evaluator LMs trained via joint training often show lower performance compared to single-format trained evaluator LMs, which indicates negative task transfer. Specifically, evaluator LMs trained only on direct assessment formats obtain higher correlations compared to jointly trained evaluator LMs across different model scales. Similarly, evaluator LMs trained only on pairwise ranking formats obtain higher average accuracy compared to multi-task trained evaluator LMs when using Mixtral-8x7B as a base model.
@@ 6.2 Is the Effectiveness of Weight Merging due to Model Ensembling?
While we empirically find that weight merging works effectively, it is unclear what might be the reason. One natural assumption might be that weight merging works effectively due to the effect of ensembling multiple models. To check the validity of this hypothesis, we conduct an ablation experiment by training multiple evaluator LMs on different random seeds and merging them. Specifically, we merge two evaluator LMs trained on direct assessment formats (denoted as ‘Direct Assessment & Direct Assessment’) and two evaluator LMs trained on pairwise ranking formats (denoted as ‘Pairwise Ranking & Pairwise Ranking’). We use Mistral-7B-Instruct as our base model.
Results are shown in Table 7. Against our expectations, we observe that in the majority of cases, merging evaluator LMs trained on the same evaluation format does not improve evaluation performances. Specifically, on direct assessment benchmarks, merging two evaluator LMs trained on direct assessment harms performance on average. Similarly, on pairwise ranking benchmarks, merging two evaluator LMs trained on pairwise ranking also harms performance on average. In contrast, by merging two evaluator LMs each trained on direct assessment and pairwise ranking formats, the resulting evaluator LM shows superior performance compared to different settings. This suggests that the positive task transfer that occurs from weight merging comes from unifying different evaluation formats, not by ensembling multiple models.
Figure 3: Finding the Optimal Alpha Value Direct Assessment performances (colored in green) and Pairwise Ranking performances (colored in blue) when altering the $\alpha$ value to merge evaluator LMs trained on different formats.
@@ 6.3 Quantifying Positive Transfer across Evaluation Formats
To explore how training on direct assessment feedback data influences pairwise ranking accuracy and vice versa, we experiment by adjusting the $\alpha$ value during linear merging. We evaluate the average performance using all eight benchmarks in our experiments. To illustrate the average performance (colored in black), we adjust the scale by multiplying direct assessment Pearson correlations, originally from 0 to 1, by 100 before averaging with pairwise ranking accuracy.
The results are shown in Figure 3. For direct assessment benchmarks, evaluator LMs obtain the optimal performance when $\alpha$ is set to 0.5. This indirectly indicates that both pairwise ranking and direct assessment feedback data contribute equally. On the other hand, for pairwise ranking benchmarks, the performance is optimal when $\alpha$ is set to 0.3. This also indirectly implies that while both benefit each other, training on pairwise ranking improves direct assessment performance more significantly than the reverse.
@@ 7 Conclusion
We introduce Prometheus 2, an open-source language model specialized in evaluating other responses. Unlike existing open evaluator language models that cannot effectively process both direct assessment and pairwise ranking—the two most prevalent evaluation schemes— the Prometheus 2 models demonstrate superior performance and consistent results on both schemes, significantly narrowing the gap with proprietary LM-based evaluations. To train the Prometheus 2 models, we develop the Preference Collection, the first pairwise ranking dataset that includes over 1,000 instance-wise evaluation criteria beyond basic qualities such as helpfulness and harmlessness. Notably, we find that merging evaluator LMs trained on either direct assessment or pairwise ranking formats can lead to a unified evaluator LM with strong performance. We hope that our work encourages more research on using open-source language models as evaluators, moving away from reliance on proprietary models for fair and accessible evaluations.
@@ Acknowledgements
We thank Sungdong Kim, Seonghyeon Ye, Sohee Yang, Dongkeun Yoon, and Hyeonbin Hwang for their helpful comments and discussions.
@@ Appendix A Quality Verification of the Preference Collection
To ensure the quality of the Preference Collection, particularly the generated verbal feedback ${v}_{{r}_{m},{r}_{n}}$, we employ five annotators with backgrounds in natural language processing. We randomly sample 200 instances with different instructions and conduct a three-part verification process. First, we assess the coherence of ${v}_{{r}_{m},{r}_{n}}$ with the scoring decision (i.e., ’A is better’ or ’B is better’). Second, we evaluate the suitability of ${v}_{{r}_{m},{r}_{n}}$ against the evaluation criteria $e$. Lastly, to determine the criticality of the feedback, we compare the newly generated ${v}_{{r}_{m},{r}_{n}}$ with a concatenation of ${v}_{{r}_{m}}$ and ${v}_{{r}_{n}}$. This aims to determine if ${v}_{{r}_{m},{r}_{n}}$ effectively leverages the mutual information between ${r}_{m}$ and ${r}_{n}$. Annotators then vote on whether ${v}_{{r}_{m},{r}_{n}}$ or the concatenation of ${r}_{m}$ and ${r}_{n}$ is more critical. The results are shown in Table 8.
Appendix B Training and Inference Hyperparameters
The configurations we used for prompting and training evaluator LMs are shown in Table 9, 10, 11. For Auto-J, PairRM and UltraRM, we utilize their prompt template, inference hyperparameter mentioned in the model cards or github repositories in order to ensure the configuration is optimal for a fair performance comparison. For proprietary LMs, Prometheus 1, and Prometheus 2 models, we use the same prompt template and evaluation configurations.
@@ Appendix C Direct Assessment Results: Extended
Table 12 and 13 (on the next page) shows the extended results Table 3. Even when changing the metrics to either Kendall-Tau and Spearman, the overall trends are maintained. Prometheus 2 shows superior evaluation performances among the open evaluator LMs, achieving high correlations with humans and proprietary LMs.
Appendix D Consistency Experiment Results: Extended
We test if evaluator LMs could give consistent scoring decisions in direct assessment formats. We inferencing multiple times with non-deterministic decoding (e.g., using temperature 1.0). Following the experimental design from Ye et al. (2023), we choose to inference 3 times and report the Krippendorff’s alpha value. As shown in Table 14, the results indicate that training on feedback data only slightly improves consistency. On the other hand, we find that the LMs with a large number of parameters achieve high consistency. This indicates the importance of selecting a large LM as the base model when training an evaluator LM. Notably, Prometheus-2-8x7B obtains the highest correlation among open evaluator LMs.
@@ Appendix D Consistency Experiment Results: Extended
Vicuna Ben. MT Ben. FLASK HHH Align. MT Ben. H.J.
Mistral-Instruct-7B
Linear Merging 0.642 0.543 0.645 78.73 67.25
DARE Merging 0.534 0.567 0.584 78.28 67.75
Mixtral-Instruct-8x7B
DARE Merging 0.685 0.665 0.659 85.52 71.96
Table 16: Pearson correlations between evaluator LMs merged with different merging methods and GPT-4-1106. Evaluator LMs trained with weight merging outperform single-format-trained and joint-trained evaluator LMs across multiple benchmarks.
Appendix E Merging Method Ablation
In Table 16, we try different merging methods introduced in our previous section. We empirically find that merging evaluator LMs with Task Arithmetic (Ilharco et al., 2022) and TIES merging (Yadav et al., 2024) constantly results in a model that degenerates. On the other hand, for the Mistral-7B based evaluator LMs, we find that linear merging and DARE merging (Yu et al., 2023) results in a model that does not degenerate and could process both evaluation formats. Also, for Mixtral-8x7B based evaluator LMs, we find that only DARE merging works effectively for both base models.
Appendix F Preference Collection Augmentation Prompt

### v2

@@ 6.1 Weight Merging vs Joint Training
Surprisingly, evaluator LMs trained via joint training often show lower performance compared to those trained only in single-format, which indicates negative task transfer. Specifically, evaluator LMs trained only on direct assessment formats obtain higher correlations compared to their jointly trained counterparts across different model scales. Similarly, evaluator LMs trained solely on pairwise ranking formats achieve higher average accuracy compared to those trained on multiple tasks, particularly when using Mixtral-8x7B as the base model.
@@
Figure 3: When merging models, the influence of relative evaluation on absolute evaluation is greater than the influence of absolute evaluation on relative evaluation. Performance of Direct Assessment (colored in green) and Pairwise Ranking (colored in blue) when altering the $\alpha$ value to merge evaluator LMs trained on different formats.
@@ 6.2 Is the Effectiveness of Weight Merging due to Model Ensembling?
While we empirically find that weight merging is effective, the underlying reason remains unclear. A natural assumption is that this effectiveness results from the ensembling effect of combining multiple models. To test this hypothesis, we conduct an ablation experiment where we train multiple evaluator LMs on different random seeds and merge them. Specifically, we merge two evaluator LMs trained on direct assessment formats (denoted as ‘Direct Assessment & Direct Assessment’) and two evaluator LMs trained on pairwise ranking formats (denoted as ‘Pairwise Ranking & Pairwise Ranking’). We use Mistral-7B-Instruct as our base model.
The results are presented in Table 6. Across multiple benchmarks, merging evaluator LMs trained on the same evaluation format does not enhance evaluation performance. Specifically, merging two evaluator LMs trained on the same evaluation format—whether direct assessment or pairwise ranking—negatively impacts performance on average for both direct assessment and pairwise ranking benchmarks. In contrast, merging two evaluator LMs, each trained on direct assessment and pairwise ranking formats, results in superior performance compared to the other settings. This indicates that the beneficial task transfer in weight merging arises from integrating different evaluation formats, not ensembling multiple models.
@@ 6.3 Quantifying Positive Transfer across Evaluation Formats
To explore how training on direct assessment feedback data influences pairwise ranking accuracy and vice versa, we experiment by adjusting the $\alpha$ value during linear merging. We evaluate the average performance using all eight benchmarks in our experiments. To illustrate the average performance (colored in black), we adjust the scale by multiplying the Pearson correlations from direct assessment, which originally range from 0 to 1, by 100 before averaging them with the pairwise ranking accuracy.
The results are shown in Figure 3. For direct assessment benchmarks, evaluator LMs obtain the optimal performance when $\alpha$ is set to 0.5. This indirectly indicates that both pairwise ranking and direct assessment feedback data contribute equally. On the other hand, for pairwise ranking benchmarks, the performance is optimal when $\alpha$ is set to 0.3. This also implies that while both benefit each other, training on pairwise ranking improves direct assessment performance more than the reverse.
@@ 7 Conclusion
We introduce Prometheus 2, an open-source LM specialized in evaluating other responses. Unlike existing open evaluator LMs that cannot effectively process both direct assessment and pairwise ranking—the two most prevalent evaluation schemes— the Prometheus 2 models demonstrate superior performance on both schemes, significantly narrowing the gap with proprietary LM-based evaluations. To train the Prometheus 2 models, we develop the Preference Collection, the first pairwise ranking dataset that includes over 1,000 instance-wise evaluation criteria beyond basic qualities such as helpfulness and harmlessness. Notably, we find that merging evaluator LMs trained on either direct assessment or pairwise ranking formats can lead to a unified evaluator LM with strong performance. We hope that our work encourages more research on using open-source LMs as evaluators.
@@ Acknowledgements
We thank the KAIST AI LKLab members for helpful discussions. This work was partly supported by LG AI Research grant (Self-improving logical reasoning capabilities of LLMs, 2024, 50%) and the Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2024-00397966, Development of a Cybersecurity Specialized RAG-based sLLM Model for Suppressing Gen-AI Malfunctions and Construction of a Publicly Demonstration Platform, 50%).
Limitations
Evaluation is fundamentally a very multi-faceted task. In this paper, we used an indirect method to assess the evaluation capability of evaluator LMs by measuring if they perform evaluations similar to human evaluators or proprietary LMs, such as GPT-4-1106 and Claude-3-Opus. However, this may not necessarily be the best approach. Future work could explore meta-evaluation pipelines that reevaluate the results of evaluator LMs or methodologies that allow humans to efficiently review evaluation results. Also note that it is crucial to use model-based evaluations in conjunction with human evaluation instead of solely relying on it.
Additionally, the degree to which evaluator LMs can generalize was based on an analysis by Kim et al. (2023), which checked for overlap between the data used to train the evaluator LMs and the data used to evaluate them. This study extended the evaluation to eight different datasets with human judgments to check the generalization capability of evaluation under various circumstances. However, this may not be sufficient. One of the major challenges in evaluating evaluator LMs is obtaining the “evaluation results” (e.g., human judgment). Automating evaluations with LMs could greatly benefit many areas of NLP research, hence the role of future work in creating feedback benchmarks that include human judgment or data for training evaluator LMs is crucial.
One downside of the Prometheus 2 is that it operates only on a 1-5 point Likert scale for absolute evaluation or a comparative evaluation style of ‘A is better & B is better’. Depending on the use case, people may need a 1-10 point absolute evaluation, a ranking method for five responses at once, or a checklist-based evaluation not covered in the paper. While proprietary LMs can flexibly conduct evaluations in any format if a well-described prompt is devised, open-source LMs cannot produce good evaluation results without training, and conversely, if trained in one or two formats, they lose the flexibility to conduct different evaluations. Future work could examine whether evaluator LMs trained in each format, as done in this paper, can handle evaluations for added formats well when weight merging is employed.
Lastly, the paper presents an evaluation model that can handle both absolute and comparative evaluation formats well through weight merging based on empirical experiments. However, fundamentally explaining why weight merging works well remains a challenging task. To address this, Section 6 indirectly analyzes the effectiveness of weight merging by comparing it with joint training, demonstrating that the improvement in evaluation performance is not due to model ensembling, and showing that the impact of comparative evaluation on absolute evaluation is greater than the reverse. Our best current interpretation is that "absolute and comparative evaluations are not completely different tasks, so weight merging could handle both without degeneration, and conversely, because they are not too similar, weight merging performed better than joint training." Future work could theoretically analyze this or further explore whether weight merging can effectively work in fields other than LLM evaluation.
@@ Appendix A Quality Verification of the Preference Collection
To ensure the quality of the Preference Collection, particularly the generated verbal feedback ${v}_{{r}_{m},{r}_{n}}$, we employ five annotators with backgrounds in natural language processing. The annotation study was designed and administered in accordance with [Affiliation X]’s ethical guidelines. Crowd workers were informed of the potential risks of participation and researcher contact information before hand in the study consent form. The hourly wage and expected study time were informed in the Prolific platform. We compensated workers 9 GBP per hour. 3 were from USA and 2 were from Asian demographics.
We randomly sample 200 instances with different instructions and conduct a three-part verification process. First, we assess the coherence of ${v}_{{r}_{m},{r}_{n}}$ with the scoring decision (i.e., ’A is better’ or ’B is better’). Second, we evaluate the suitability of ${v}_{{r}_{m},{r}_{n}}$ against the evaluation criteria $e$. Lastly, to determine the criticality of the feedback, we compare the newly generated ${v}_{{r}_{m},{r}_{n}}$ with a concatenation of ${v}_{{r}_{m}}$ and ${v}_{{r}_{n}}$. This aims to determine if ${v}_{{r}_{m},{r}_{n}}$ effectively leverages the mutual information between ${r}_{m}$ and ${r}_{n}$. Annotators then vote on whether ${v}_{{r}_{m},{r}_{n}}$ or the concatenation of ${r}_{m}$ and ${r}_{n}$ is more critical. The results are shown in Table 7. Note that the Preference Collection only includes English instances.
Appendix B Training and Inference Details
The configurations we used for prompting and training evaluator LMs are shown in Table 8, 9, 10. For Auto-J, PairRM and UltraRM, we utilize their prompt template, inference hyperparameter mentioned in the model cards or github repositories in order to ensure the configuration is optimal for a fair performance comparison. For proprietary LMs, Prometheus 1, and Prometheus 2 models, we use the same prompt template and evaluation configurations.
For both training and inference, we utilized eight 40GB NVIDIA A100 GPUs. Training required approximately 800 GPU hours, using the implementation from the Alignment Handbook repository [^2] https://github.com/huggingface/alignment-handbook. For inference, we used the vllm framework [^3] https://github.com/vllm-project/vllm.
The results from Direct Assessment are averaged after three multiple runs, and pairwise grading is conducted in a single run. Instead of using error bars, we report the consistency in assessment formats, Krippendorff’s alpha for consistency in direct assessment, and transitivity statistics for consistency in pairwise ranking.
@@ Appendix C Direct Assessment Results: Extended
Table 11 and 12 (on the next page) shows the extended results Table 3. Even when changing the metrics to either Kendall-Tau and Spearman, the overall trends are maintained. Prometheus 2 shows superior evaluation performances among the open evaluator LMs, achieving high correlations with humans and proprietary LMs.
Appendix D License
Our models are released under the Apache 2.0 license. The Preference Collection dataset is subject to OpenAI’s Terms of Use for generated data. The model could be used for commercial purposes while the dataset is intended for research purposes. We used perspective API to ensure that the training data or evaluation datasets do not include PII-included instances.
Evaluator LM Vicuna Bench MT Bench FLASK Feedback Bench
Llama2-Chat 7B 0.183 0.203 0.065 0.070 0.229 0.186 0.211 0.419
Llama2-Chat 13B 0.145 0.146 -0.019 0.037 0.160 0.174 0.174 0.453
Llama2-Chat 70B 0.282 0.382 0.150 0.196 0.310 0.310 0.231 0.487
Mistral-Instruct-7B 0.314 0.391 0.208 0.281 0.395 0.384 0.287 0.454
Mixtral-Instruct-8x7B 0.395 0.468 0.433 0.419 0.410 0.408 0.304 0.551
Prometheus-7B 0.405 0.425 0.290 0.263 0.282 0.251 0.236 0.770
Prometheus-13B 0.397 0.434 0.299 0.352 0.365 0.352 0.299 0.793
Auto-J (13B) 0.282 0.242 0.303 0.272 0.312 0.282 0.312 0.515
Prometheus-2-7B 0.543 0.476 0.390 0.372 0.476 0.446 0.377 0.784
Prometheus-2-8x7B 0.559 0.515 0.535 0.483 0.526 0.507 0.388 0.800
GPT-3.5-Turbo-0613 0.255 0.287 0.148 0.157 0.360 0.315 0.298 0.489
GPT-4-1106 / 0.553 / 0.590 / 0.609 0.517 0.662
Claude-3-Opus 0.553 / 0.590 / 0.609 / 0.453 0.693
Table 11: Kendall-Tau correlations between reference evaluators (listed on top) and evaluator LMs. The best comparable statistics are bolded and second best underlined except proprietary LMs.
Evaluator LM Vicuna Bench MT Bench FLASK Feedback Bench
Llama2-Chat 7B 0.236 0.255 0.084 0.089 0.301 0.244 0.279 0.511
Llama2-Chat 13B 0.178 0.179 -0.025 0.044 0.206 0.222 0.224 0.543
Llama2-Chat 70B 0.348 0.466 0.197 0.252 0.391 0.389 0.298 0.585
Mistral-Instruct-7B 0.389 0.480 0.266 0.358 0.499 0.478 0.374 0.563
Mixtral-Instruct-8x7B 0.476 0.556 0.545 0.517 0.505 0.500 0.386 0.659
Prometheus-7B 0.508 0.528 0.385 0.349 0.367 0.326 0.317 0.876
Prometheus-13B 0.492 0.534 0.401 0.470 0.474 0.454 0.398 0.893
Auto-J (13B) 0.337 0.297 0.408 0.365 0.402 0.358 0.408 0.623
Prometheus-2-7B 0.664 0.591 0.509 0.482 0.597 0.555 0.491 0.885
Prometheus-2-8x7B 0.660 0.615 0.669 0.605 0.642 0.618 0.496 0.912
GPT-3.5-Turbo-0613 0.319 0.354 0.192 0.198 0.446 0.390 0.374 0.565
GPT-4-1106 / 0.659 / 0.721 / 0.729 0.650 0.753
Claude-3-Opus 0.659 / 0.721 / 0.729 / 0.567 0.784
Table 12: Spearman correlations between reference evaluators (listed on top) and evaluator LMs. The best comparable statistics are bolded and second best underlined except proprietary LMs.
Evaluator LM HHH Alignment MT Bench Human Judg. Auto-J Eval
Auto-J (13B) 46.61 75.57 28.96 48.14 69.12 20.98 47.40 76.64 29.24
Prometheus-2-7B 74.21 74.66 0.45 63.24 70.78 7.54 68.11 75.07 6.96
Claude-3-Opus 84.62 94.57 9.95 62.65 77.65 15.00 61.04 82.90 21.86
Evaluator LM Vicuna Ben. MT Ben. FLASK
Llama2-Chat 7B 0.3558 0.2565 0.4379
Llama2-Chat 13B 0.2017 0.2998 0.4038
Llama2-Chat 70B 0.5212 0.4559 0.6204
Mistral-Instruct-7B 0.5157 0.4393 0.5884
Mixtral-Instruct-8x7B 0.5459 0.6229 0.6976
Prometheus-7B 0.6049 0.5363 0.5970
Prometheus-13B 0.5734 0.5181 0.5624
Auto-J (13B) 0.4976 0.5069 0.6160
Prometheus-2-7B 0.6018 0.5340 0.5991
Prometheus-2-8x7B 0.6383 0.6862 0.7874
GPT-3.5-Turbo-0613 0.7108 0.4800 0.6389
GPT-4-1106-preview 0.7366 0.8271 0.8355
Claude-3-Opus 0.8284 0.8601 0.8976
Table 14: Krippendorff’s alpha statistics for evaluator LMs when prompted 3 times via non-deterministic decoding.
Evaluator LM Preference Collection
Transitivity
Mistral-Instruct-7B 87.10
Mixtral-Instruct-8x7B 90.45
Pair RM 91.40
Ultra RM 94.25
Auto-J (13B) 89.65
Prometheus-2-7B 97.60
Prometheus-2-8x7B 96.75
GPT-3.5-Turbo-0613 84.35
GPT-4-1106-preview 95.70
Claude-3-Opus 96.20
Table 15: Transitivity statistics to measure consistency in pairwise ranking evaluation settings.
Appendix E Consistency of Evaluator LMs
In addition to obtaining high correlation and accuracy, achieving high consistency is another important aspect for evaluator LMs. We first test if evaluator LMs could give consistent scoring decisions in direct assessment formats. We inferencing multiple times with non-deterministic decoding (e.g., using temperature 1.0). Following the experimental design from Ye et al. (2023), we choose to inference 3 times and report the Krippendorff’s alpha value. As shown in Table 14, the results indicate that training on feedback data only slightly improves consistency. On the other hand, we find that the LMs with a large number of parameters achieve high consistency. This indicates the importance of selecting a large LM as the base model when training an evaluator LM. Notably, Prometheus-2-8x7B obtains the highest correlation among open evaluator LMs.
@@ Appendix E Consistency of Evaluator LMs
Lastly, we conduct an experiment to test if evaluator LMs could achieve consistent scores across different evaluation formats. To do this, we use pairwise ranking benchmarks and measure the performance differences when prompted with direct assessment formats and pairwise ranking formats. Specifically, following Kim et al. (2023), to process pairwise ranking datasets in a direct assessment scheme, we evaluate each response separately and compare the scoring decisions. We mark it as correct if the evaluator LM provides a higher score for the human-chosen response over the rejected one. As shown in Table 13 (on the previous page), the results show that Prometheus 2 models show lower performance differences across evaluation formats, indicating their robustness.
Evaluator LM BiGGen Bench FLASK
Reference-free Reference-based $\Delta$ Reference-free Reference-based $\Delta$
Mistral-Instruct 0.305 0.310 0.005 0.331 0.374 0.043
Mixtral-Instruct 0.320 0.322 0.002 0.377 0.386 0.009
Prometheus-2-7B 0.403 0.455 0.052 0.425 0.545 0.120
Prometheus-2-8x7B 0.424 0.472 0.048 0.411 0.555 0.144
GPT-3.5-Turbo-0613 0.236 0.252 0.016 0.354 0.374 0.020
GPT-4-1106 0.554 0.599 0.045 0.616 0.679 0.063
Table 16: Pearson correlations between different evaluator models with and without the reference answer and Human. Reference-based evaluations outperform reference-free evaluations across all evaluator LMs.
Merging Method Direct Assessment Benchmarks Pairwise Ranking Benchmarks Average
Vicuna Ben. MT Ben. FLASK (Human) Feedback Ben. Average HHH Align. MT Ben. H.J. Auto-J Pref. Ben. Average
Linear 0.642 0.543 0.544 0.878 0.652 78.73 67.25 73.80 92.45 78.06 82.93
Slerp 0.648 0.532 0.536 0.879 0.649 74.66 70.2 72.33 92.60 77.44 82.67
Task Arithmetic 0.518 0.497 0.482 0.831 0.582 80.09 69.80 72.82 93.00 78.93 81.01
TIES 0.534 0.567 0.529 0.826 0.614 79.64 67.75 72.91 93.95 78.56 80.58
Dare-Ties 0.653 0.545 0.543 0.880 0.655 79.64 66.57 74.68 93.30 78.55 83.27
Dare-Linear 0.666 0.548 0.545 0.882 0.660 74.66 70.78 75.07 93.25 78.44 83.32
Table 17: Pearson correlations and accuracy measurements across various benchmarks for different merging methods. The best comparable statistics are bolded and second best underlined.
Appendix F Reference-free Evaluation in Direct Assessment Formats
In this section, we assess the impact of excluding a reference answer in evaluations conducted using direct assessment formats. The results are presented in Table 16 (on the previous page). For this experiment, we employ FLASK (Ye et al., 2023) which includes human judgments and additionally the BiGGen Bench (Kim et al., 2024). The BiGGen Bench is a generation benchmark which includes a evaluation criteria tailored to each instance and provides 2840 human judgments (excluding the multilingual tasks) in direct assessment formats.
Across both benchmarks and different evaluator LM variants, the correlation with humans diminishes when the reference answer is discarded. Even for GPT-4-1106, there is a significant performance degradation (0.045, 0.063). This suggests that including a reference answer is crucial for conducting effective evaluations with LMs. Interestingly, Prometheus-2-7B achieves better performance in a reference-free setting (0.403, 0.425) than Mistral-7B-Instruct-v0.2 (0.310, 0.374). Similar trends are observed for Prometheus-2-8x7B (0.424, 0.411) and Mixtral-8x7B-Instruct-v0.1 (0.322, 0.386). This implies that one effect of training an evaluator LM with a reference answer included is to induce the ability to ground judgments to the given reference answer.
Appendix G Merging Method Ablation
In this section, in addition to linear merging, we also test different merging techniques including:
  • Slerp merging (Goddard et al., 2024) operates by interpolating two weights ${\theta}_{d}$ and ${\theta}_{p}$ while preserving the geometric properties of the spherical space in which ${\theta}_{d}$ and ${\theta}_{p}$ reside. Specifically, this is conducted by normalizing ${\theta}_{d}$ and ${\theta}_{p}$ into unit length and then merging the two weights based on the coefficient $\alpha$ such as:
    ${\theta}_{final}=\alpha\times\frac{{\theta}_{d}}{||{\theta}_{d}||}+(1-\alpha)\times\frac{{\theta}_{p}}{||{\theta}_{p}||}$ (4)
    $(1-\alpha)\times({\theta}_{p}-{\theta}_{init})$
  • DARE merging (Yu et al., 2023), while also similar to Task Arithmetic and TIES merging, performs a Random Drop and Re-scale operations in the task vector ${\theta}_{d}-{\theta}_{init}$ and ${\theta}_{p}-{\theta}_{init}$ to remove redundant weights. We find that DARE merging work best when we choose Mixtral-8x7B as our base model. DARE-linear merging is what was originally proposed by Yu et al. (2023). In DARE-TIES merging, the Elect operation from Yadav et al. (2024) is additionally added after the Re-scale operation.
We conduct our experiments based on the implementation from MergeKit (Goddard et al., 2024) [^4] https://github.com/arcee-ai/mergekit
In Table 17 (on the previous page), we measure the performance of evaluator LMs employing different merging methods. In direct assessment benchmarks, DARE-Linear achieves the best performance, followed by DARE-TIES and Linear merging. In pairwise ranking benchmarks, Task Arithmetics achieves the best performance, with only a minimal difference compared to other methods. On average, DARE-Linear performs best. Based on these results, we have trained Prometheus-2-7B with DARE-Linear merging. We also opted to train Prometheus-2-8x7B using DARE-Linear merging. Although the optimal merging method might differ, we have not conducted additional experiments due to computational limitations. Future work could explore whether these findings hold true.
Appendix H Preference Collection Augmentation Prompt